OpenStudy (usukidoll):

dy/dx = x+1 / y separable equation

OpenStudy (usukidoll):

\[\frac{dy}{dx} = \frac{x+1}{y}\]

OpenStudy (usukidoll):

\[y dy = (x+1) dx\]

OpenStudy (usukidoll):

\[\frac{1}{2} y^2 = \frac{1}{2}x^2+x+C\]

OpenStudy (anonymous):

That is correct!

OpenStudy (usukidoll):

\[y^2 = x^2+2x+2C\]

OpenStudy (anonymous):

correct...

OpenStudy (usukidoll):

\[y = \sqrt{x^2+2x+2c}\] with the plus minus stuff can't do it in latex

OpenStudy (rational):

C is as good an arbitrary constant as 2C

hartnn (hartnn):

`\(\pm\)` \(\pm\)